OpenStudy (anonymous):

solve, write answer in set builder notation. x-9 > -7 (x+2) > is underlined

OpenStudy (anonymous):

\[x-9\geq-7x-14\] \[x+7x\geq -14+9\] \[8x\geq -5\] \[x\geq -\frac{5}{8}\]

OpenStudy (anonymous):

set builder notation \[\{x|x\geq -\frac{5}{8}\}\]

OpenStudy (anonymous):

thank you.... I have another ? listed and no response yet.... it is about 15-20 down the list or so if you want to give it a shot.

OpenStudy (anonymous):

ok let me find it